OpenStudy (anonymous):

A container was pressurized with 150.0 kPa. What was the initial volume if the final volume is 20.4 L with a pressure of 40.0 kPa if there were no temperature change? I'm pretty sure it's Boyle's Law P1V1=P2V2

OpenStudy (anonymous):

How do you solve for V1?

thomaster (thomaster):

\(\Large P_1V_1=P_2V_2\) \(\Large 150*V_1=40*20.4\) \(\Large V_1=\dfrac{P_2V_2}{P_1}\) \(\Large V_1=\dfrac{40*20.4}{150}\)

OpenStudy (anonymous):

Thanks!